# Generative diffusion model surrogates for mechanistic agent-based biological models

### v2

@@ 1 Introduction
Applying neural networks as surrogates for the CPM presents additional challenges compared to that of deterministic systems of PDEs such as heat transfer. Agent-based models often describe emergent behaviors that arise from the interaction of multiple agents within the model that are not explicitly encoded beforehand. CPM models are stochastic, making deterministic neural-network architectures ill-posed to serve as model surrogates. Despite these challenges, the authors have developed a neural-network based surrogate for a CPM of vasculogenesis, where biological vascular patterns emerge from the actions of individual cell agents reacting to each other and a chemotactic diffusive field governed by a system of PDEs [19]. This surrogate predicted model configurations at an intermediate timestep of 100 Monte-Carlo steps for a single set of parameters, and was unable to replicate model configurations at long timescales or for multiple areas of a parameter space. Because of the stochastic nature of the CPM model, model configurations at long timescales do not retain memory of a reference configuration, or the configuration many computational timesteps prior to the current configuration. Therefore, two equivalent samples from a given area of a model’s parameter space may have different explicit configurations while demonstrating the same phenotypic features that reflect the emergent behaviors of the model mechanisms. These are additional challenges towards developing a surrogate model for a CPM.
@@ 2.2 Cellular-Potts Model simulation and training data generation
Approximately 1000 CPM cell agents were placed randomly throughout the 256x256 simulation domain at the beginning of the simulation and settled into vascular-like structures over the first 10,000 CPM simulation timesteps, referred to as Monte-Carlo steps (MCS) of the simulation. This simulation state was saved as an initial timepoint $t_{0}$ to generate training data for the DDPM surrogate. Training data for the surrogate model was created by saving binary images of the cell layer of model configurations from the CPM model. Images representing model configurations were saved every 100 MCS for a given simulation between 15000 and 20,001 MCS. This procedure was repeated for 100 unique simulations per parameter set spanning a 2-dimensional parameter space. Unique parameter sets were created by defining the parameter $J_{\text{cell,medium}}$ as values within the set $\{0,5,10,15,20\}$. We also defined the diffusive field decay constant parameter $k$ to be $\{0.05,0.1875,0.325,0.465,0.6\}$ to define a 5x5 grid of unique parameter sets (Figure 1A). Over the course of each simulation, the initial configuration $t_{0}$ with the baseline parameters defined in Table 1 evolved to a new configuration representative of the new parameter set (Figure 1B, Movie 1). This yielded 5100 images per area of the parameter space and a total of 127500 images for all 25 classes for the total dataset. This dataset was then used to train the diffusion model to serve as a surrogate. This same dataset was also used to train the image classifier for surrogate model verification. A separate dataset was generated in the same manner consisting of 510 images per area of the parameter space for a total of 12750 images used for evaluation of the trained classifier.
@@ 2.3 Denoising Diffusion Model configuration and training
To create our CPM surrogate, we used the modified DDPM model architecture, training schema, and fast deterministic sampling method as previously described in the EDM2 method [26] to create a class-conditional DDPM. We use the edm2-img64-s preset as defined in the original publication but modify the model inputs to allow for single-channel images as well as vary EDM2 training parameters P_mean and P_std to define a more optimal set for our use case as is recommended in the original EDM2 publication. P_mean corresponds to an average amount of gaussian noise to be applied to an image during training of the DDPM, and P_std corresponds to the variability of noise about P_mean. We use this model preset to perform diffusion directly in pixel space of our single-channel 256x256 images for a total of 12,748,800 training steps. We chose this setup because at the time of this investigation, the training presets described produce state-of-the-art DDPM model performance for image synthesis for both generation in pixel space and latent space of a pretrained autoencoder for natural images. Thus, we expect these procedures to allow for good performance for our intended use case. Training was performed on 16 V100 GPUs.
@@ 3.1 Surrogate Model Selection
Figure 3: Image classifier identifies mode collapse in trained candidate DDPM surrogate models. DDPM surrogate model performance is sensitive to training hyperparameters P_mean and P_std. Each candidate model was used to generate 20,000 configurations uniformly distributed among the 25 classes. Generated outputs were evaluated for uniformity with a trained image classifier.
@@
We performed a coarse-grained study of training parameters P_mean and P_std to attempt to identify a set of parameters that would best allow the DDPM surrogate learn the distributions of model configurations defined in the training set. We defined P_mean to range from -1.6 to 1.6, and P_std to range from 1.0-3.0 for a total of 25 candidate models. This parameter range encompassed the parameters found to be optimal for image synthesis in pixel space when trained on natural images as described in the EDM2 work [26]. For each candidate model, we used the model snapshot saved at the final training step to generate 20,000 CPM configurations uniformly distributed among all 25 parameter spaces (classes). We evaluated each model’s ability to generate CPM configurations within all classes by verifying generated outputs with the trained EfficientNetV2-s classifier. For each trained DDPM surrogate candidate, we plotted the frequency CPM configurations generated in the set of 20,000 CPM configurations (Figure 3). Evaluating the candidate surrogate models in this manner revealed that certain combinations of P_mean and P_std lead to mode collapse, or a loss in the diversity of output of the DDPM model. Of the candidate models, the surrogate trained with P_mean = 0.533 and P_std = 2.333 was selected because generated samples were roughly uniform. This model was then used to generate model configurations unsupervised in application as the CPM surrogate model.
@@ 4 Discussion
Neural network-based surrogates allow for use of the graphics processing unit (GPU) to accelerate computational speed of a given task. The Cellular-Potts models as implemented in the widely available open-source frameworks CompuCell3D [1], Morpheus [2], Chaste [29], and Artistoo [3] to name a few, evaluate the Cellular-Potts algorithm on the central processing unit (CPU). Additionally, these frameworks must explicitly evaluate model configurations at each MCS which results in an increased computational requirement and prolonged simulation evaluation time. Investigation into GPU-accelerated implementations for the Cellular-Potts algorithm is an active field of research. Recent methods allow for parallel evaluation of the Cellular-Potts algorithm using the GPU which allows for significant speed increases for model evaluation [30, 31]. However, these GPU-accelerated Cellular-Potts implementations are not compatible with the widely adopted open-source CPM frameworks and may have limited ability to accommodate multi-method simulation such as diffusive fields described by systems of PDEs. DDPMs and other neural network-based surrogates may be an additional method for allowing for GPU acceleration of mechanistic agent-based models [19].
@@
We have demonstrated the novel use of image classifiers as a verification technique to assist in trained DDPM model selection and to evaluate a model’s potential for performance. In contrast, the Frechet’s Inception Distance (FID) [32] is a metric developed for evaluating the generative modeling performance of generative adversarial networks (GANs) and has become popular for the evaluation of DDPMs as well [26, 33, 34]. This scalar value defines a similarity of two distributions of images which is useful for comparing generative models trained on similar datasets such as the widely used open-source natural image dataset ImageNet [35]. However, a scalar value is ill-suited to confirm for a user whether an image or generated model configuration exists within a specific parameter space or class. A lower FID would increase the likelihood that a generated sample exists within the intended class, but it is not confirmation of whether a specific generated sample does exist within the intended class. An image classifier, however, can confirm this. Additionally, the FID may be less effective for our use case as many areas of the CPM parameter space yield visually similar configurations that may be difficult to identify in isolation. A trained image classifier will be able to confirm the identity of a generated configuration.
@@
A limitation of our DDPM model was the inability to achieve a perfectly uniform distribution of generated samples as made obvious by the classifier verification. This phenomena is an example of mode collapse and is a common occurrence in generative models [36]. Mode collapse becomes apparent when a generative model is unable to demonstrate the diversity of data present in the training set. Mode collapse is a common occurrence in the training of generative models. In addition to model selection, the trained image classifier could have been used in combination with the DDPM to ensure that all generations of the surrogate are of the intended class. This would define a hybrid DDPM-classifier surrogate whereby the DDPM could be sampled repeatedly until a sample of the intended class is generated. This would allow the use of a classifier to “rescue” a DDPM that may be difficult to optimally train at the cost of computational time and efficiency during generation. A hybrid surrogate model using the classifier together with the DDPM would address the observed mode collapse. Other proposed methods to alleviate mode collapse for generative diffusion models include elaborate architectures such as Diffusion-GAN hybrid models [37] or creative regularization and reinforcement learning-based fine-tuning strategies [38]. Diffusion models may also be controlled to produce outputs of interest via model guidance. Classifier guidance has been popularized by Dhariwal and Nichol [39] whereby gradients of a pre-trained classifier steer sample generation from a DDPM towards a particular class or desired output. More recently, a method of classifier-free guidance has been described that requires the combination of a trained conditional and unconditional diffusion model to control generative sample diversity [40]. Finally, a method of diffusion model autoguidance has been described that allows for the control and improvement of a DDPM generative output using a less-trained version of itself [34]. Future efforts will investigate the potential for these techniques to improve upon a DDPM surrogate for a CPM.
@@
At long timescales of 20,000 MCS, memory of an initial or reference configuration is lost. However, certain characteristics of the configuration such as the number and approximate size of circular vascular lacunae, the width of vascular branches, and the number of vascular sprouts are all functions of the underlying mechanistic model parameter values. These characteristic metrics define the phenotype or state of the biological system at a given set of parameters and were investigated in the original publication that described the CPM [4]. Additionally, the stochastic nature of these CPMs give rise to differences in model configurations between two simulation replicates with identical underlying parameters and initial configurations. Therefore, the characteristics of the class of configurations that arise due to a given set of parameters may be of primary importance to the investigation and application of stochastic agent-based mechanistic biological models. The image classifier that we have defined in this work has potential to define these characteristics and encode the universality class of model configurations. This has significant implications for stochastic multicellular agent-based models as this now enables model interoperability between equivalent models encoded in different frameworks. At present, minor differences in algorithmic implementation of the for model methods such as the CPM result in drastic numerical differences for equivalent models encoded in different open-source frameworks of the same underlying mathematical method. This has been cited as a common and significant barrier to the interoperability, reusability, and reproducibility of stochastic multicellular models in systems biology [41].
@@
While we use DDPMs in this work to model stochastic biological systems, these generative diffusion models are being adapted to study other stochastic physical systems. DDPMs and other generative model architectures have been used as model surrogates for the simulation of calorimeter showers to investigate fundamental particle experiments [21, 42]. Additionally, DDPMs are a key component in CorrDiff, cited as a component in a digital twin for weather forecasting [24]. This work adapting DDPMs as surrogate models for multicellular biological systems is a step towards diffusion model-based biological and medical digital twins.
In conclusion, we have leveraged generative diffusion models as effective surrogates for stochastic, multicellular, mechanistic biological models. We additionally described the training and application of an image classifier in defining the universality classes of equivalent stochastic simulation configurations. Finally, we use this classifier to select and verify optimal DDPM-based surrogate models. Our surrogate yields representative simulation states at long timescales in advance of a reference configuration and achieves a  22x speed increase as compared to native code execution. Our work represents a step towards generative diffusion model based digital twins for biological and medical systems.

### v3

@@ 1 Introduction
Applying neural networks as surrogates for the CPM presents additional challenges compared to that of deterministic systems of PDEs such as heat transfer. Agent-based models often describe emergent behaviors that arise from the interaction of multiple agents within the model that are not explicitly encoded beforehand. CPM models are stochastic, making deterministic neural-network architectures ill-posed to serve as model surrogates. Despite these challenges, the authors have developed a neural-network based surrogate for a CPM of vasculogenesis, where biological vascular patterns emerge from the actions of individual cell agents reacting to each other and a chemotactic diffusive field governed by a system of PDEs [19]. This surrogate predicted model configurations at an intermediate timestep of 100 Monte-Carlo steps for a single set of parameters, and was unable to replicate model configurations at long timescales (see Figure S1 and Movie S1) or for multiple areas of a parameter space. Because of the stochastic nature of the CPM model, model configurations at long timescales do not retain memory of a reference configuration, or the configuration many computational timesteps prior to the current configuration. Therefore, two equivalent samples from a given area of a model’s parameter space may have different explicit configurations while demonstrating the same phenotypic features that reflect the emergent behaviors of the model mechanisms. These are additional challenges towards developing a surrogate model for a CPM.
@@ 2.2 Cellular-Potts Model simulation and training data generation
Approximately 1000 CPM cell agents were placed randomly throughout the 256x256 simulation domain at the beginning of the simulation and settled into vascular-like structures over the first 10,000 CPM simulation timesteps, referred to as Monte-Carlo steps (MCS) of the simulation. This simulation state was saved as an initial timepoint $t_{0}$ to generate training data for the DDPM surrogate. Training data for the surrogate model was created by saving binary images of the cell layer of model configurations from the CPM model. Images representing model configurations were saved every 100 MCS for a given simulation between 15000 and 20,001 MCS. This procedure was repeated for 100 unique simulations per parameter set spanning a 2-dimensional parameter space. Unique parameter sets were created by defining the parameter $J_{\text{cell,medium}}$ as values within the set $\{0,5,10,15,20\}$. We also defined the diffusive field decay constant parameter $k$ to be $\{0.05,0.1875,0.325,0.465,0.6\}$ to define a 5x5 grid of unique parameter sets (Figure 1A). This sampling of the parameter space was selected to provide a variety of visually distinct phenotypes. While many of these may be less biologically relevant to the developing vasculature that the original CPM was developed to represent, we expected that this variety of visually apparent phenotypes may pose a more challenging dataset for a surrogate to replicate. We also intended for this to result in a method that is generalizable to other computational models, as different visually apparent phenotypes often represent differing system states that would be appropriate to capture with a surrogate model. Over the course of each simulation, the initial configuration $t_{0}$ with the baseline parameters defined in Table 1 evolved to a new configuration representative of the new parameter set (Figure 1B, Movie 1). This yielded 5100 images per area of the parameter space and a total of 127500 images for all 25 classes for the total dataset. This dataset was then used to train the diffusion model to serve as a surrogate. This same dataset was also used to train the image classifier for surrogate model verification. A separate dataset was generated in the same manner consisting of 510 images per area of the parameter space for a total of 12750 images used for evaluation of the trained classifier.
@@ 2.3 Denoising Diffusion Model configuration and training
To create our CPM surrogate, we used the modified DDPM model architecture, training schema, and fast deterministic sampling method as previously described in the EDM2 method [26] to create a class-conditional DDPM. We use the edm2-img64-s preset as defined in the original publication but modify the model inputs to allow for single-channel images as well as vary EDM2 training parameters P_mean and P_std to define a more optimal set for our use case as is recommended in the original EDM2 publication. P_mean corresponds to an average amount of gaussian noise to be applied to an image during training of the DDPM, and P_std corresponds to the variability of noise about P_mean. We use this model preset to perform diffusion directly in pixel space of our single-channel 256x256 images for a total of 12,748,800 training steps. We chose this setup because at the time of this investigation, the training presets described produce state-of-the-art DDPM model performance for image synthesis for both generation in pixel space and latent space of a pretrained autoencoder for natural images. Thus, we expect these procedures to allow for good performance for our intended use case. Training was performed on 16 V100 GPUs. This preset had the additional benefit of having the smallest computational footprint of the defined presets and thus may be easier to train and more accessible to the broader scientific community.
@@ 3.1 Surrogate Model Selection
Figure 3: Image classifier identifies mode collapse in trained candidate DDPM surrogate models. DDPM surrogate model performance is sensitive to training hyperparameters P_mean and P_std. Each candidate model was used to generate 20,000 configurations uniformly distributed among the 25 classes, or discrete areas of the CPM parameter space. (A) Generated outputs were classified using a trained image classifier and sample counts per class are displayed as a heatmap in the equivalent 5x5 grid layout shown previously in Figure 1B. Sample counts per class (or discrete parameter set) are displayed within each corresponding cell of each heatmap subplot. KL divergence between generated sample distributions and a uniform distribution is displayed for each DDPM model to provide a quantitative measure of uniformity. KL divergence values close to 0 indicate a more uniform distribution among classes in generated samples. (B) Confusion matrix plots demonstrating classifier identified class (detected class) compared with the conditioning label for generation (intended class) for each model. Generation accuracy is displayed for each DDPM model.
@@
We performed a coarse-grained study of training parameters P_mean and P_std to attempt to identify a set of parameters that would best allow the DDPM surrogate learn the distributions of model configurations defined in the training set. We defined P_mean to range from -1.6 to 1.6, and P_std to range from 1.0-3.0 for a total of 25 candidate models. This parameter range encompassed the parameters found to be optimal for image synthesis in pixel space when trained on natural images as described in the EDM2 work [26]. For each candidate model, we used the model snapshot saved at the final training step to generate 20,000 CPM configurations uniformly distributed among all 25 parameter spaces (classes). Each sample was generated with a unique random seed (0-19,999 for a total of 20,000 unique samples). We evaluated each model’s ability to generate CPM configurations within all classes by verifying generated outputs with the trained EfficientNetV2-s classifier. Generated class distributions were evaluated for uniformity through calculating the KL divergence between the class-distribution of generated samples and a uniform distribution for each candidate DDPM model. For each trained DDPM surrogate candidate, we plotted the class frequency of CPM configurations generated from the set of 20,000 (Figure 3A). We additionally plot confusion matrices to determine the accuracy of class-conditional generation (intended class) as compared with the sample that was generated as classified by the classifier (detected class) in Figure 3B. Evaluating the candidate surrogate models in this manner revealed that certain combinations of P_mean and P_std lead to mode collapse, or a loss in the diversity of output of the DDPM model. Candidate models were expected to produce samples uniformly distributed across the 25 classes because the training data consisted of a uniform distribution among the classes. Of the candidate models, the surrogate trained with P_mean = 1.6 and P_std = 1.667 was selected because it generated the most uniform distribution of samples by measure of KL divergence from a uniform distribution, and the greatest class accuracy from class-conditional generation. This model surrogate was then used to generate mechanistic model sample configurations in an unsupervised manner in application as the CPM surrogate model.
We also quantified the vascular morphologic features of branch width and distributions of lacunae area for generated samples. We used Earth Mover’s Distance (EMD) to quantify distances between distributions of these morphologic features from DDPM generated configurations and the training data for each class, for each model (Figure S2). Large EMD values indicate a distribution of features for samples of a specific class that deviate significantly from the corresponding training data for that class or area of the CPM parameter space. Absent values indicate samples from which these features could not be quantified, such as samples consisting entirely of background with no generated vessel network.
@@ 4 Discussion
Neural network-based surrogates allow for use of the graphics processing unit (GPU) to accelerate computational speed of a given task. The Cellular-Potts models as implemented in the widely available open-source frameworks CompuCell3D [1], Morpheus [2], Chaste [29], and Artistoo [3] to name a few, evaluate the Cellular-Potts algorithm on the central processing unit (CPU). Additionally, these frameworks must explicitly evaluate model configurations at each MCS which results in an increased computational requirement and prolonged simulation evaluation time. Investigation into GPU-accelerated implementations for the Cellular-Potts algorithm is an active field of research. Recent methods allow for parallel evaluation of the Cellular-Potts algorithm using the GPU which allows for significant speed increases for model evaluation [30, 31]. However, these GPU-accelerated Cellular-Potts implementations are not compatible with the widely adopted open-source CPM frameworks and may have limited ability to accommodate multi-method simulation such as diffusive fields described by systems of PDEs. Because of this, we were unable to perform a hardware-fair timing comparison between the DDPM surrogate on the GPU, and the Cellular-Potts model implementation on the GPU. Given the multiple iterative denoising steps required for sample generation from a DDPM, CPU evaluation of the DDPM surrogate would be computationally costly and disadvantageous for the purposes of a model surrogate. DDPMs and other neural network-based surrogates may be an additional method for allowing for GPU acceleration of mechanistic agent-based models [19].
@@
We have demonstrated the novel use of image classifiers as a verification technique to assist in trained DDPM model selection and to evaluate a model’s potential for performance. In contrast, the Frechet’s Inception Distance (FID) [32] is a metric developed for evaluating the generative modeling performance of generative adversarial networks (GANs) and has become popular for the evaluation of DDPMs as well [26, 33, 34]. This scalar value defines a similarity of two distributions of images which is useful for comparing generative models trained on similar datasets such as the widely used open-source natural image dataset ImageNet [35]. However, a scalar value is ill-suited to confirm for a user whether an image or generated model configuration exists within a specific parameter space or class. A lower FID would increase the probability that a generated sample exists within the intended class, but it is not confirmation of whether a specific generated sample does exist within the intended class. The FID is a global similarity metric calculated on the scale of tens-of-thousands of image samples rather than a per-sample confirmation of similarity. An image classifier, however, can confirm this similarity on a per-sample basis. Additionally, the FID may be less effective for our use case as many areas of the CPM parameter space yield visually similar configurations that may be difficult to identify in isolation. A trained image classifier will be able to confirm the identity of a generated configuration.
At present, our surrogate generates representative configurations at long timescales in 25 discrete areas of the parameter space. Generating reference configurations between these discrete parameter sets would be useful, however between-class interpolation to achieve this off-grid generalization is a complicated problem. The model described in this work is a class-conditional DDPM and methods of effectively interpolating between class labels is an area of active research for DDPM models. Multiple recent publications propose methods to achieve such interpolation including training additional neural network architectures to add conditioning controls to DDPM generation [36], interpolating between generated samples in the latent space of a latent DDPM [37], or mixing score functions of conditional generations [38]. While these methods prove useful for controlling generation of natural images, controlled generation between classes is particularly complicated in our use-case as a mechanistic model surrogate. Generating between-class samples derived through methods that rely on the discretely sampled training data does not guarantee that the resulting between-class image is representative of the corresponding area of the parameter space created by the mechanistic model. Additional methods must be developed to validate that between-class samples generated by the DDPM map effectively to those of the original mechanistic model. This is a non-trivial research activity that will be the focus of ongoing and future work.
An additional limitation of our DDPM model was the inability to achieve a perfectly uniform distribution of generated samples in some hyperparameter configurations as made obvious by the classifier verification and KL divergence. This phenomenon is an example of mode collapse and is a common occurrence in generative models [39]. Mode collapse is the phenomenon when a generative model is unable to demonstrate the diversity of data present in the training set. This is apparent in Figure 3 for DDPM configurations that fail to generate samples in all classes, or uniformly distributed across classes, even though the training data was uniformly distributed.
@@
In addition to model selection, the trained image classifier could have been used in combination with the DDPM to ensure that all generations of the surrogate are of the intended class. To do this, the DDPM surrogate could be repeatedly sampled until the classifier confirms that a sample from the intended class was generated. This configuration would define a hybrid DDPM-classifier surrogate where the classifier component may “rescue” a DDPM that may be difficult to optimally train at the cost of computational time and efficiency during generation. A hybrid surrogate model using the classifier together with the DDPM would address the observed mode collapse. Other proposed methods to alleviate mode collapse for generative diffusion models include elaborate architectures such as Diffusion-GAN hybrid models [40] or creative regularization and reinforcement learning-based fine-tuning strategies [41]. Diffusion models may also be controlled to produce outputs of interest via model guidance. Classifier guidance has been popularized by Dhariwal and Nichol [42] whereby gradients of a pre-trained classifier steer sample generation from a DDPM towards a particular class or desired output. More recently, a method of classifier-free guidance has been described that requires the combination of a trained conditional and unconditional diffusion model to control generative sample diversity [43]. Finally, a method of diffusion model autoguidance has been described that allows for the control and improvement of a DDPM generative output using a less-trained version of itself [34]. Future efforts will investigate the potential for these techniques to improve upon a DDPM surrogate for a CPM.
@@
At long timescales of 20,000 MCS, memory of an initial or reference configuration is lost. However, certain characteristics of the configuration such as the number and approximate size of circular vascular lacunae, the width of vascular branches, and the number of vascular sprouts are all functions of the underlying mechanistic model parameter values. These characteristic metrics define the phenotype or state of the biological system at a given set of parameters and were investigated in the original publication that described the CPM [4]. Additionally, the stochastic nature of these CPMs give rise to differences in model configurations between two simulation replicates with identical underlying parameters and initial configurations. Therefore, the characteristics of the class of configurations that arise due to a given set of parameters may be of primary importance to the investigation and application of stochastic agent-based mechanistic biological models. The image classifier that we have defined in this work has potential to define these characteristics and encode the universality class of model configurations. This has significant implications for stochastic multicellular agent-based models as building upon this in future work may enable model interoperability between equivalent models encoded in different frameworks. At present, minor differences in algorithmic implementation of the for model methods such as the CPM result in drastic numerical differences for equivalent models encoded in different open-source frameworks of the same underlying mathematical method. This has been cited as a common and significant barrier to the interoperability, reusability, and reproducibility of stochastic multicellular models in systems biology [44].
@@
While we use DDPMs in this work to model stochastic biological systems, these generative diffusion models are being adapted to study other stochastic physical systems. DDPMs and other generative model architectures have been used as model surrogates for the simulation of calorimeter showers to investigate fundamental particle experiments [21, 45]. Additionally, DDPMs are a key component in CorrDiff, cited as a component in a digital twin for weather forecasting [24].
Our work demonstrates that a conditional DDPM architecture and its training procedure previously defined for generating natural images may be applied as a surrogate for Cellular-Potts models of biological systems. Our surrogate currently performs forward inference of 20,000 MCS across a discretely sampled parameter space. Looking ahead, an ideal digital twin would be capable of performing this inference at parameter sets between the demonstrated 25 classes, have some capability to solve for the inverse problem of parameter inference for mapping to its real-world biological counterpart, and allow for on-line updating. We previously discussed adapting data-driven between-class interpolation methods to provide a practical step towards delivering fine-grained forward inference across the parameter space[36, 37, 38]. For the inverse problem, we may consider adapting image-based parameter estimation methods that currently exist for mapping computational models to images of their biological counterparts[46] as part of a greater digital twin infrastructure. For on-line updating, recent work leveraging diffusion models as aligned diffusion schrodinger bridges [47] allows for learning the stochastic dynamics in between samples such that conditional generation of a resulting configuration given an arbitrary reference may occur. Thus, incorporating these existing advancements onto our work defines one potential path towards diffusion-model based digital twins.